If z= x+ i y ,
p(z^z,z^z) = p(2z,z+i) , find z ?
(p) Means permutation

Answer by ikleyn(52788)

.
If z= x+ i y ,
p(z^z,z^z) = p(2z,z+i) , find z ?
(p) Means permutation
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


The meaning of this post is as dark as a midnight in a forest at no moon.